A garden has 60 feet of bordering and the length of the garden is to be 2 feet greater than the width. Find the length and the width of the garden

Answer by checkley79(3341)

P=2L+2W
60=2*(W+2)+2W
60=2W+4+2W
4W=60-4
4W=56
W=56/4
W=14 ANS/ FOR THE WIDTH.
L=14+2=16 ANS. FOR THE LENGTH.
PROOF:
60=2*16+2*14
60=32+28
60=60